You are here

Projects Updates for Geothermal Monitoring Well on Bardeen Quad


Search tips:

  • This form will search for words in the title OR the description. If you would like to search for the same term(s) across both the title and description, enter the same search term(s) in both fields.
  • This form will search for any of the words you enter in a field, not the exact phrase you enter. If you would like to search for an exact phrase, put double quotes (") around the phrase. For example, if you search for Bike Path you will get results containing either the word Bike OR the word Path, but if you search for "Bike Path" you will get results containing the exact phrase Bike Path.
  1. Geothermal monitoring borehole on the Bardeen Quad

    Hi everyone,

    I wanted to report that yesterday we completed the thermal response test in the borehole, and our work at the site is done. Currently, we are working on processing and analyzing the data. This work should be completed later this week, and when done we will share the results with Sachin and the rest of the project team.

    Tim Stark will also use these results to compare with the wellbore model his student is developing.





    Andrew Stumpf, Ph.D, LG, P.Geo

    Associate Geologist

    Illinois State Geological Survey

    Prairie Research Institute

    University of Illinois at Urbana-Champaign

    615 East Peabody Drive, Champaign, Illinois USA 61820

  2. Fiber-optic cable intact

    Hi everyone,

    I can report today that the fiber-optic cable is intact and the test returned a signal along the entire 235 m cable. What you see in the photograph is data along the cable going down and coming back up (470 m total).

    We have ordered the drop box to store the fiber-optic cable, and a hole has been dug for it. The box will be installed by F&S after the holidays. The fencing is being left around the site until this is done.

    Wishing you all the best this holiday season.


    Attached Files: 
  3. Borehole sealed

    Hi Everyone, We successfully installed the geothermal loop and fiber optic cable today. The borehole is now grouted closed. My appreciation goes out to the group of colleagues and staff from the ISGS that assisted the drill crew to insert the 400-feet of tubing in the ground.

    The drill crew will be back in the morning to clean up the site and move equipment offsite. We still need to install a drop box to store the fiber-optic cable and access the geothermal loop. The box has been ordered and if it arrives in the morning, we may be able to get it installed before the holidays. In about a month we will start measuring the ground temperature and run a DTRT test once the grout cures.

    Happy Holidays!!


  4. Installing fiber optics

    Hi everyone, We had a successful day of collecting geophysical logs in the borehole. I have seen print outs of this excellent data, and is an important part of the geologic characterization. Thanks goes out to Tim Young for coming out of retirement to help us complete the work.

    Tomorrow we will install the U-bend and fiber-optic cable in the borehole. Our task is to sink a 400-foot log tube filled with water (weighing ~600 lbs) down the hole. For those nearby you may want to drop by for a visit. Once this is done, we will proceed to grout the borehole closed.



  5. Last drilling update

    Hi everyone,

    We finished drilling today and advanced the borehole to a depth of 385 feet.

    Tomorrow we will collect geophysical logs of the borehole. If you want to drop by we should be in full swing by 11 am. It will take some time to clean out any material that has collapsed in the borehole overnight and make sure the hole is full of mud before logging the open hole.



  6. Daily update

    Hi everyone, As expected our rate of drilling today slowed when drilling the shale bedrock. The crew was able to drill an additional 25 feet to a depth of 355 feet.

    Next week, since Monday will be the last day we have to drill the hole deeper, we have decided to stop coring and just drill straight through so that we can reach the target depth of 450 feet. Tuesday is needed to clean out the borehole so that geophysical logging can be done on Wednesday. On Thursday and Friday, we will install the geothermal loop and fiber optic cable and grout the hole closed.

    Have a great weekend.


  7. Daily update

    Hi Everyone, We are starting out Wednesday morning at 294 feet, about 5 feet into the bedrock. The crew will be widening the hole this morning from 5 inches to 5.9 inches before going deeper into bedrock. Looks like a nice day for drilling!!


  8. Daily update

    Hi everyone,

    After a morning of thawing out the drilling system from well below freezing temperatures on the weekend, we were able to reach 240 feet. With the top bedrock being <60 feet, it is likely they will reach it on Tuesday.

    My colleague Jason Thomason will be at the site on Tuesday while I am out of town.



  9. Daily update

    Hi everyone,

    We had a successful day 2 after deicing the system. We are now down to 165 feet and heading into the oldest glacial strata. If you are following along with the log I sent, we are about 170 feet on it.

    We will start up again Monday morning.



  10. daily update

    Hi everyone,


    After day 1 we have drilled to a depth of 90 feet. This is excellent progress with snow much of the day, and not getting in the ground until late morning.

    So far, the geology is similar to previous drilling in the area (see log attached).




  11. Funding Approval for Geothermal Characterization/Monitoring Station

    Evan De Lucia and Mohamed Attalla approved $65,610 of funding from the Carbon Credit Sales Fund for a Geothermal Characterization/Monitoring Station at John Bardeen Quad.


    An email of approval is attached below.

    The project proposal is attached below.

Subscribe to